OpenStudy (anonymous):

fan and medal one easy question For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x - 2y = -4 3x + y = 12 a.-2y - 4 b.2y - 4 c.2y + 4 d.-2y + 4

OpenStudy (anonymous):

@amistre64 do u think u can help?

OpenStudy (amistre64):

just isolate x in the first eqauton

OpenStudy (anonymous):

by isolate do you mean take away or like put in (x) @amistre64

OpenStudy (amistre64):

get the little x all by itself

OpenStudy (anonymous):

do i just add x to the - x @amistre64

OpenStudy (amistre64):

add x to both sides is fine -x - 2y = -4 +x +x ------------ -2y = x-4 what next

OpenStudy (anonymous):

plus 4 on each side, right? -2y+4=x ? @amistre64

OpenStudy (amistre64):

very good

OpenStudy (anonymous):

thats all!?

OpenStudy (amistre64):

well, the next step is to pick the option that matches it ...

OpenStudy (anonymous):

of course! So D right? haha

OpenStudy (amistre64):

yep :)

OpenStudy (anonymous):

Thank you soooo much :)